# WildFireCan-MMD: A Multimodal Dataset for Classification of User-Generated Content During Wildfires in Canada

### v3

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. In this work, we focus on multimodal wildfire social media data, which, although existing in current datasets, is currently underrepresented in Canadian contexts. We present WildFireCan-MMD, a new multimodal dataset of X posts from recent Canadian wildfires, annotated across twelve key themes. We evaluate zero-shot vision-language models on this dataset and compare their results with those of custom-trained and baseline classifiers. We show that while baseline methods and zero-shot prompting offer quick deployment, custom-trained models outperform them when labelled data is available. Our best-performing custom model reaches 84.48±0.69% f-score, outperforming VLMs and baseline classifiers. We also demonstrate how this model can be used to uncover trends during wildfires, through the collection and analysis of a large unlabeled dataset. Our dataset facilitates future research in wildfire response, and our findings highlight the importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@
Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,MarziehAmini@cunet.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca
@@ 1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi (2023); Kumar et al. (2024b), mental health monitoring Ghanadian et al. (2024); Dalal et al. (2024), detecting and combating hate speech in digital environments Kiritchenko et al. (2021); Kumar et al. (2024a), pandemic discourse analysis Imran et al. (2022) and sentiment and behavioral trend analysis Ma et al. (2024); Curto et al. (2024). Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam et al. (2018); Ofli et al. (2020); Alam et al. (2021); Dwarakanath et al. (2021); Algiriyage et al. (2022); Hossain et al. (2022); Khattar and Quadri (2022); Basit et al. (2023); Koshy and Elango (2023); Rezk et al. (2023). Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information. The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston et al. (2015). This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop an automated method for sifting through this information.
To foster research in this area, Alam et al. (2018, 2021) developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising 1589 samples, and is specific to the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset Alam et al. (2021) are limited to a single month in 2016, as well as another month of California samples from 2018. Therefore, wildfires in Canada remain underexplored in social media analysis research.
@@
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several Vision-Language Models (VLMs), custom deep-learning classifiers, as well as classical approaches, as a minimum achievable performance. Our best-performing classifier utilizes RoBERTa Liu (2019) and Vision Transformer (ViT) models to jointly process the text and image components of tweets. A custom transformer-based fusion module is designed to enable the modalities to interact deeply. While WildFireCan-MMD is sourced from data posted between 2022 and 2024, we collected an additional set of unlabeled tweets spanning seven years (2018-2024). Using our best-performing classifier, we analyze this unlabeled data and uncover trends and patterns of social media use during wildfires in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available at [Anonymized GitHub], under CC-BY License. The contributions of this study are summarized as:
@@ 2.1 Social media analysis for disaster response
Significant research has been conducted on the role of social media during disasters. Fauzi (2023) conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions. Acikara et al. (2023) highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information. Muniz-Rodriguez et al. (2020) focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. In other works, Dwarakanath et al. (2021), Bukar et al. (2020), and Mundottukandi et al. (2024) discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage et al. (2022) discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios.
Our work is motivated by this literature as well as the literature shown in Appendix G which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
@@ 2.2 Multimodal classification for disaster response
Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters Olteanu et al. (2014); Dong et al. (2021); Ochoa and Comes (2021), it is now generally considered that multimodal approaches offer stronger results Ofli et al. (2020); Abavisani et al. (2020); Ramirez et al. (2025).
@@
The study by Alam et al. (2018) released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes). CrisisMMD contains thousands of tweets labeled for categories like “infrastructure and utility damage” or “rescue volunteering or donation effort”.
@@
Additionally, Alam et al. (2018) later released the HumAID Alam et al. (2021) dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019). It is similarly annotated for emergency response categories (e.g. casualties, infrastructure damage).
@@
There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex Olteanu et al. (2015) and CrisisMMD Alam et al. (2018) datasets Wu et al. (2022). CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Ofli et al. (2020); Khattar and Quadri (2022); Basit et al. (2023); Koshy and Elango (2023); Rezk et al. (2023) and many others. Therefore, there is a notable gap in the in the availability of multimodal wildfire datasets.
@@
In the context of Canada, where the most common and destructive natural disaster is wildfires, CrisisMMD is lacking in data. Although it contains  18k samples from other disasters such as hurricanes, earthquakes and floods, the amount of wildfire data is proportionally small (1589 samples). Additionally, this data is only from the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016, as well as another month of California samples from 2018. Therefore, wildfires in Canada remain underexplored in social media analysis research.
To address this limitation, this work focuses on gathering social media data related to Canadian wildfires, and proposes WildFireCan-MMD, a new Canadian wildfire social-media dataset with a wildfire-specific taxonomy 12 categories.
@@
Algorithms: Previous work by Ofli et al. (2020) introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures respectively, followed by a dense layer for prediction. The work by Basit et al. (2023), expands on Ofli et al. (2020) by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT Lan et al. (2020) as a text embedder, and DeIT Touvron et al. (2021) as an image embedder, provided the best results for the eight-class classification task. Koshy and Elango (2023) implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa Liu (2019) model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam et al. (2018).
@@
Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix G. Most previous studies use a CNN for the image modality and then a BERT-based transformer for the text modality. Other works sometimes use ViT for the image modality. Concatenative fusion of the two modalities into a series of linear layers seems to be the most common approach, with other works using multiplicative or additive fusion. Cross-attention fusion of the two modalities seems to be another popular approach that seems to give promising results. The approach given by Pranesh (2022) proposes a transformer-like structure to fuse the two modalities. We build on the this previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
@@
In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Zhang et al. (2024); Li et al. (2022). Specifically, we evaluate three VLMs, GPT-4 Achiam et al. (2023), LLaVA Liu et al. (2023), and SmolVLM Marafioti et al. (2025), to assess performance across a range of capabilities and computational requirements. GPT-4 serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
@@ 3.1 Data Collection
To construct the dataset, X’s Pro-Tier paid API was utilized to collect a diverse set of posts, including images and text, surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from the wildfire season (April-September) of each respective year. Data collection was conducted using a hashtag-based search strategy as seen in other works Alam et al. (2018); Gupta and Hewett (2020); Cruickshank and Carley (2020). Appendix A presents the full queries used for data collection. To curate the hashtags, basic hashtags such as "#wildfire, #forestfire" were first used. Then, through an iterative process of adding new hashtags by collecting sample tweets, more relevant hashtags were discovered. This approach of using hashtags was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, hashtags identified through iterative experimentation were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region from which the tweet originated.
3.2 Taxonomy Development
To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst (2022). Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic. The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15. Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords. Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics. Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased. For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected. Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear. Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders. Less useful posts were also noticed, such as advertisements and political debates/complaints. Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix C.
@@ 3.2 Taxonomy Development
3.3 Data Labeling
Labeling was carried out in-house by three independent annotators using Label Studio [^2] https://labelstud.io/, using the defined taxonomy presented in Table 1. Examples of labeled tweets are presented in Figure 1.
Each image-text pair was given a single label. In cases where the label could differ depending on whether text or image was being considered, preference was given to the modality that provided greater value for that particular sample. For example, in Figure 1(c) the image might be labeled ‘Smoke & Air Quality’, but the text might be ‘Warnings & Status Updates’. In this case, ‘Warnings & Status Updates’ would be chosen as the label as the text contains more valuable information than the image alone. If neither contained detailed information about the time or location like the text does in this example, then the sample would be labeled as ‘Smoke & Air Quality’ as that is the only value that would otherwise be provided. The size of labeled classes is presented in Table 1.
Annotation was done by three independent annotators, all authors of this paper. The annotators are male, 20-30 year old, fluent in English and reside in Canada. First, initial rounds of manual labeling on small subsets of the data gave a good understanding of what was being discussed and shown in the data, and confidence was gained in the decided classes. Over the course of labeling, samples were flagged that contained personal information.
A vote between the labels from the three annotators decided the final labels. Annotation agreement scores can be seen in Table 2. A majority agreement was reached on 88.4% of the labels. Taking a vote between all annotators for the final label and comparing the result with each annotator, we can see which annotator was closest to the voted labels. From the results, annotators 1 & 2 had the best agreement scores. This is further demonstrated by Cohen’s and Fleiss’ Kappa metrics, which indicate that a moderate-to-good level of agreement is achieved between the independent annotators. In cases of disagreement, the final label was chosen from the ‘expert’ annotator (Annotator 1), who had the most experience with the dataset collection and devising the taxonomy, and also had the highest agreement score with other annotators. Additional statistics can be found in Appendix B.

### v4

@@ Abstract
Rapid information access is vital during wildfires, yet traditional data sources are slow and costly. Social media offers real-time updates, but extracting relevant insights remains a challenge. In this work, we focus on multimodal wildfire social media data, which, although existing in current datasets, is currently underrepresented in Canadian contexts. We present WildFireCan-MMD, a new multimodal dataset of posts on X from recent Canadian wildfires, annotated across twelve key themes. We evaluate zero-shot vision-language models on this dataset and compare their results with those of custom-trained and baseline classifiers. We show that while baseline methods and zero-shot prompting offer quick deployment, custom-trained models outperform them when labelled data is available. Our best-performing custom model reaches 84.48±0.69% f-score, outperforming VLMs and baseline classifiers. We also demonstrate how this model can be used to uncover trends during wildfires, through the collection and analysis of a large unlabeled dataset. Our dataset facilitates future research in wildfire response, and our findings highlight the importance of tailored datasets and task-specific training. Importantly, such datasets should be localized, as disaster response requirements vary across regions and contexts.
@@
Braeden Sherritt1, Isar Nejadgholi2, Efstratios Aivaliotis1, Khaled Mslmani1, Marzieh Amini1 1Carleton University, Ottawa, Canada 2National Research Council Canada, Ottawa, Canada BraedenSherritt@cmail.carleton.ca,isar.nejadgholi@nrc-cnrc.gc.ca,Efstratiosaivaliotis@cmail.carleton.ca, KhaledMslmani@cmail.carleton.ca, MarziehAmini@cunet.carleton.ca
@@ 1 Introduction
Large-scale analysis of online interactions, using machine learning algorithms has enabled critical applications such as disaster response Fauzi2023; kumar2024extracting, mental health monitoring ghanadian2024socially; dalal2024review, detecting and combating hate speech in digital environments kiritchenko2021confronting; kumar2024detecting, pandemic discourse analysis tbcov19 and sentiment and behavioral trend analysis ma2024toward; curto2024crime. Among these applications, analysis of social media data during natural disasters has gained increasing importance in recent years Alam2018; Ofli2020; alam2021humaid; Dwarakanath2021; Algiriyage2022; Hossain2022; Khattar2022; Basit2023; Koshy2023; Rezk2023. Such data contains crucial details such as reports of casualties, infrastructure damage, and urgent needs of affected individuals, among other critical information. The effective use of the shared information can facilitate humanitarian aid efforts for more efficient management and faster recovery Houston2015. This valuable data, however, is mixed in with noise, such as humorous content, advertisements, and other irrelevant information. Therefore, there is a critical need to develop an automated method for sifting through this information.
To foster research in this area, Alam2018; alam2021humaid developed two multimodal datasets, CrisisMMD and HumAID, both collected from Twitter, including several disasters such as earthquakes and hurricanes. However, the classes of these datasets are generalized across all types of disasters, meaning they are not tailored to specific disaster scenarios. As a result, the classification scheme is broad, and certain nuances or unique information that may be relevant to a particular type of disaster could be overlooked. Also, while the CrisisMMD dataset does contain wildfire data from the California wildfires of 2017, the proportion of this type of data compared to the rest is relatively small, only comprising 1589 samples, and is specific to the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset alam2021humaid are limited to a single month in 2016, as well as another month of California samples from 2018. Therefore, wildfires in Canada remain underexplored in social media analysis research.
@@
To benchmark the available classification models on WildFireCan-MMD, we evaluated the zero-shot classification capabilities of several Vision-Language Models (VLMs), custom deep-learning classifiers, as well as classical approaches, as a minimum achievable performance. Our best-performing classifier utilizes RoBERTa roberta2019 and Vision Transformer (ViT) models to jointly process the text and image components of tweets. A custom transformer-based fusion module is designed to enable the modalities to interact deeply. While WildFireCan-MMD is sourced from data posted between 2022 and 2024, we collected an additional set of unlabeled tweets spanning seven years (2018-2024). Using our best-performing classifier, we analyze this unlabeled data and uncover trends and patterns of social media use during wildfires in Canada. The full set of labeled data (containing the tweet IDs and labels), as well as the experimental code, is available on Github (https://github.com/Multimodal-Social-Media-Data-Analysis/WildfireCanMMD). The contributions of this study are summarized as:
@@ 2.1 Social media analysis for disaster response
Significant research has been conducted on the role of social media during disasters. Fauzi2023 conducted a bibliometric analysis of social media use in disaster response and offered a comprehensive overview of key studies and emerging research directions. Acikara2023 highlighted the potential of social media analytics in disaster response by leveraging users as “citizen sensors” to rapidly gather critical, time-sensitive information. Muniz2020 focused on the public health aspects of social media use during emergencies, noting its effectiveness in spreading emergency warnings and assessing needs after a disaster. In other works, Dwarakanath2021, 9220091, and 10485423 discussed the crucial role of social media in managing crises, pointing out how platforms like Facebook and Twitter support communication and coordination among affected individuals and communities. Furthermore, in a systematic review of the application of deep learning in disaster response, Algiriyage2022 discussed the successes and challenges faced in this area, proposed guidelines for future research, and highlighted the importance of using multimodal data in disaster scenarios. Our work is motivated by this literature as well as the literature shown in Appendix F which focuses this work on using machine learning models to triage Canada-specific multimodal social media data to inform disaster response and management.
@@ 2.2 Multimodal classification for disaster response
Datasets: While many works utilize textual datasets sourced from social media to classify posts during disasters crisislex2014; Dong2021; Ochoa2021, it is now generally considered that multimodal approaches offer stronger results Ofli2020; Abavisani2020; ramirez2025.
@@
The study by Alam2018 released CrisisMMD, a collection of tweets ( 18k samples) with images from seven major natural disasters (including wildfires, earthquakes, floods, and hurricanes). CrisisMMD contains thousands of tweets labeled for categories like “infrastructure and utility damage” or “rescue volunteering or donation effort”.
@@
Additionally, alam2021humaid later released the HumAID dataset which consists of  77K human-labeled disaster tweets sampled from  24M tweets over 19 events (2016–2019). It is similarly annotated for emergency response categories (e.g. casualties, infrastructure damage).
@@
There are several other multimodal datasets that exist, but most are variations or combinations of the CrisisLex crisislext262015 and CrisisMMD datasets Xuehua2022. CrisisMMD has become a foundational resource for studying disaster-related communication patterns and the combined use of text and images. It remains the most prominent dataset in this domain, used in many recent studies Ofli2020; Khattar2022; Basit2023; Koshy2023; Rezk2023 and many others. Therefore, there is a notable gap in the in the availability of multimodal wildfire datasets.
@@
In the context of Canada, where the most common and destructive natural disaster is wildfires, CrisisMMD is lacking in data. Although it contains  18k samples from other disasters such as hurricanes, earthquakes and floods, the amount of wildfire data is proportionally small (1589 samples). Additionally, this data is only from the California wildfires of October 2017. The wildfire-related samples from Canada in the HumAID dataset come from only a single month in 2016 (2,259 samples), as well as another month of California samples from 2018 (7,444 samples). Therefore, wildfires in Canada remain underexplored in social media analysis research. To address this limitation, this work focuses on gathering social media data related to Canadian wildfires, and proposes WildFireCan-MMD, a new Canadian wildfire social-media dataset with a wildfire-specific taxonomy 12 categories.
@@
Algorithms: Previous work by Ofli2020 introduced a joint representation from two parallel deep learning architectures where one represents the text modality and the other represents the image modality. Image and text modalities are vectorized through VGG16 and Word2Vec+CNN architectures respectively, followed by a dense layer for prediction. The work by Basit2023, expands on Ofli2020 by experimenting with multiple recent transformer structures, for both the text and image modalities, and compared them against the baseline reported in the previous work. They found that using a combination of AlBERT albert2020 as a text embedder, and DeIT deit2021 as an image embedder, provided the best results for the eight-class classification task. Koshy2023 implemented further experiments with text and image transformers and showed improvements by using bi-LSTMs instead of LSTMs. Furthermore, they have noted that multiplicative fusion strategies produce better results on a classification task when combined with the previously described methods. Specifically, they demonstrated that using the fine-tuned RoBERTa roberta2019 model for text, Vision Transformer model for image, biLSTM+attention mechanism with multiplicative fusion, for prediction, achieves the best results for the eight-class classification task described by Alam2018.
@@
Many architectures have been proposed in other works for classification on CrisisMMD, as seen in Appendix F. Most previous studies use a CNN for the image modality and then a BERT-based transformer for the text modality. Other works sometimes use ViT for the image modality. Concatenative fusion of the two modalities into a series of linear layers seems to be the most common approach, with other works using multiplicative or additive fusion. Cross-attention fusion of the two modalities seems to be another popular approach that seems to give promising results. The approach given by pranesh2022 proposes a transformer-like structure to fuse the two modalities. We build on the this previous work and explore the existing structures on the newly annotated data, WildFireCan-MMD.
@@
In addition to fine-tuned structures, we explore using VLMs in a zero-shot setting Duzhen2024; Li2022. Specifically, we evaluate four VLMs, GPT-4o-mini OpenAI2023, LLaVA-v1.5-13B Liu2023, Qwen2.5-VL-7B qwen2025, and SmolVLM-2B smolvlm2025, to assess performance across a range of capabilities and computational requirements. GPT-4o-mini serves as a state-of-the-art, closed-source benchmark, while LLaVA and SmolVLM are open-source models that use visual instruction tuning and are optimized for varying levels of hardware, reflecting ongoing efforts to build accessible, high-performing multimodal systems.
@@ 3.1 Data Collection
To construct the dataset, X’s Pro-Tier paid API was utilized to collect a diverse set of posts, including images and text, surrounding the British Columbia and Alberta wildfires of 2022 and 2023, as well as the Jasper wildfires of 2024. The dataset comprises 4,688 unique image-text pairs gathered from the wildfire season (April-September) of each respective year. Data collection was conducted using a hashtag-based search strategy as seen in other works Alam2018; gupta2020; Cruickshank2020. Table 1 presents the full queries used for data collection. To curate the hashtags, basic hashtags such as "#wildfire, #forestfire" were first used. Then, through an iterative process of adding new hashtags by collecting sample tweets, more relevant hashtags were discovered. This approach of using hashtags was necessary due to the limited availability of geotagged content, as most users on the platform do not enable precise geo-location. Instead, hashtags identified through iterative experimentation were found to be effective indicators of relevant content originating from Canada during wildfire events. The specificity of these hashtags ensured that the collected tweets were not only on-topic but also predominantly from users in the targeted geographic regions. The dataset includes the timestamp of when the tweet was posted, the user-provided location, and, in some cases, the specific region from which the tweet originated.
Year Query
Table 1: The queries used for collecting the data for WildFireCan-MMD.
@@
3.2 Taxonomy Development
To examine the content of the tweets and discover the hidden latent semantics of the data, BERTopic was employed to perform multimodal unsupervised topic modeling Grootendorst2022. Text and image inputs were vectorized using the pre-trained “Clip-ViT-B-32” [^1] https://huggingface.co/sentence-transformers/clip-ViT-B-32 model as suggested by BERTopic. The default parameter settings of BERTopic were used, which reduced the dimensionality to 5, while keeping the size of the local neighborhood at 15. Stop words were removed, and both unigrams and bigrams were considered as candidates of topic keywords. Using BERTopic on the dataset, the algorithm determined the number of topics to categorize the data into, resulting in 100 fine-grained topics. Hierarchical topic reduction was applied using Top2Vec, and the number of topics was gradually decreased. For each iteration, the representative text and visual samples of topics and keywords generated by BERTopic were qualitatively inspected. Once the number of topics was reduced to around 15, it became clear which topics were the strongest in the dataset, and the humanitarian utility of some of these topics began to appear. Some examples of useful topics included posts about infrastructure, weather reports, and reports on the actions of first responders. Less useful posts were also noticed, such as advertisements and political debates/complaints. Some examples of these topics, along with representative images and corresponding keywords generated by BERTopic, are presented in Appendix B.